# Local Attention Mechanism: Boosting the Transformer Architecture for Long-Sequence Time Series Forecasting

### v2

@@ 4.3 Comparing LAM and ProbAttn within the same architecture framework
Following the above comparison, it is pertinent to consider whether the two attention mechanisms, LAM and ProbAttn (Informer’s attention mechanism) are competing on equal terms. Informer is a significantly more complex model than the vanilla transformer integrating LAM. The Informer model utilises its own positional encoding based on various time frequencies (seconds, minutes, hours, weeks, etc.) and incorporates layers such as one-dimensional convolutions to facilitate feature extraction from the dataset. Therefore, before proceeding with further experimentation, it is essential to compare LAM and ProbAttn by aligning the frameworks under which both attention mechanisms are implemented. To this end, we will compare the results obtained by ProbAttn and LAM within the vanilla transformer architecture, to isolate the effects of Informer’s attention mechanism with respect to the characteristics of its architecture. This involves using only fully connected layers and the sine-based positional encoding from the original proposal by Vaswani et al [20], instead of the complex architecture of Informer.
@@
The results presented in Table 4 provide a clear illustration of the performance differences. Notably, LAM outperforms ProbAttn in 37 instances, while ProbAttn prevails in 23. However, it is crucial to also examine the cumulative error across all datasets and scenarios to establish a comprehensive metric. Specifically, the cumulative MSE for LAM is $28.979$, compared to $34.841$ for ProbAttn. This indicates that, overall, LAM exhibits a lower cumulative error than ProbAttn. Similarly, for MAE, LAM achieves an accumulated value of $19.976$, whereas ProbAttn records $21.666$. Additionally, it is noteworthy that, on average, LAM reduces the error of ProbAttn by approximately 25% for longer prediction horizons. This finding corroborates the improvement and stability of our model’s results under the LSTF problem.
(a) MSE Bayesian test
(b) MAE Bayesian test
Figure 6: Bayesian tests ProbAttn vs LAM
To corroborate the data presented in the table, we include the corresponding Bayesian tests, as illustrated in Figure 6. These tests reveal that the differences are significant in favour of our proposal, with a probability of 97.2% for MSE and a probability of 93.9% for MAE of LAM being better than ProbAttn. These results support the effectiveness of our proposal within the same comparison framework.
In view of the results obtained in this experimentation, we argue that the complex architecture used in Informer is not properly justified. Therefore, the question of determining a transformer architecture that successfully exploits the structure of long-term time series forecasting problems remains open. We note that in the paper that introduced Informer [22] there is no comparison to the vanilla transformer and, thus, no empirical justification of the changes introduced at the architecture level – the study focuses on comparing ProbAttn with FullAttention within their own architecture and comparing Informer to other models specifically designed for time series forecasting. For completeness, we also include comparisons between LAM and FullAttention in Section 5.
@@ 4.4 On the suitability of benchmark models for the LSTF problem
As a conclusion to this experimentation, we have identified several areas for improvement in the literature. Firstly, all datasets utilised have a relatively small number of samples; none exceeds 100,000 instances (Table 1). Transformer models for long-sequence time series forecasting, being highly complex architectures in terms of the number of weights, require a substantial amount of data to effectively train all parameters. The current datasets are insufficient in terms of sample size for this task, which produces results that are not sufficiently conclusive. Additionally, the datasets do not always encompass a large set of features. An increase in the number of variables or features inherently increases the complexity of multivariate prediction for LSTF, leading to more interesting problems.
@@
For these reasons, we believe that the currently used datasets, while useful for measuring model performance, are insufficient to establish a comprehensive benchmark for time series analysis in general, or for the Long Sequence Time-series Forecasting (LSTF) problem in particular. Consequently, we recognise the necessity of obtaining new datasets that more accurately represent the LSTF problem to better evaluate model performance. This need will be addressed in the following section of this paper.
@@ 5.1 Experimental Details
Datasets: For this experiment, we selected four datasets: Individual Household Electric Power (IHEP) [44], New York Taxi (NYTaxi) [45], Residential Power Battery (RPB) [46], and Time-series Industrial Anomaly (TiNA) [47, 48]. The IHEP dataset contains over 200,000 samples with 9 features, representing household electricity consumption measurements over a 4-year period. The NYTaxi dataset includes over 2 million instances with 6 features, such as distance travelled, number of passengers, amount paid, and trip duration. These metrics are aggregated by summing all trips within the same hour, allowing the dataset to support predictions for subsequent hours, days, or weeks. The RPB dataset comprises more than 500,000 samples with 6 features, measuring the actual consumption of a population, solar power generation, and battery levels. Finally, the TiNA dataset includes over 38 million instances with 108 features, sampled every two seconds. This dataset is a real-world dataset provided by ArcelorMittal, reflecting data from an ore mining machine.
@@
Algorithms: In the experimentation provided in Section 4, Informer and LAM are the closest models in terms of performance. The other algorithms, including transformers such as LogTrans and Reformer, showed significantly poorer results. Consequently, we have decided to exclude these algorithms from this comparison to obtain concise and more interpretable results. Therefore, we have retained Informer and our proposed model, and added the vanilla transformer with its attention mechanism (FullAttention) to serve as a baseline. For our proposed model and the vanilla attention mechanism, the optimal depth was determined through hyperparameter tuning. The transformer architecture employed is a fully connected, symmetric structure with an equal number of encoding and decoding layers. The depth was varied between 2, 3, 5, 7, and 10 layers, contingent upon memory capacity, to identify the best configuration. This parameter will be named as number of encoding-decoding layers for the rest of the section. By using the same architecture to baseline the vanilla attention mechanism, we ensure that the observed results can be solely attributed to our proposal and the network depth. The same standardisation and metrics introduced in the previous section are used for the present one. The train/validation/test split utilised in this study allocated 70% of the data for training, with 15% of this subset further designated for validation and the remaining 85% for training. The remaining 30% of the data was used as the test set.
@@ 5.2 Results and Analysis
Table 6: Multivariate forecasting over the Individual Household Electric Power dataset (IHEP), New York Taxi dataset (NYTaxi), Residential Power Battery dataset (RPB) and Time-series Industrial Anomaly dataset (TiNA). MSE and MAE are presented in the table, the lower the better, best results in bold. Empty cells are due to the impossibility of executing the algorithm for these parameters due to memory problems given the size of the model. Layers represent the best configuration of encoding-decoding layers used to obtain the presented results. Count stands for best result counts. Mean Accumulated is computed as the added error divided by the number of valid rows in the table (non-empty). The Vanilla algorithm is the original transformer proposed in [20].
@@
As observed in the Table 7, LAM consistently emerges as the superior algorithm in this comparison. Specifically, it outperforms the other algorithms 31 times, whereas Informer obtains better results 5 times, and Vanilla wins 4 times. This demonstrates the clear superiority of our proposal over both the baseline and Informer models. Regarding the mean of the MSE and MAE across rows with valid values, Vanilla appears to have a slight edge. However, it is important to note that these results are significantly influenced by Vanilla’s inability to run under numerous circumstances due to its computational demands. Consequently, these findings strongly establish the validity of our proposal, even when utilising a fixed architecture.
@@
Table 7: Multivariate forecasting over the Individual Household Electric Power dataset (IHEP), New York Taxi dataset (NYTaxi), Residential Power Battery dataset (RPB) and Time-series Industrial Anomaly dataset (TiNA). MSE and MAE are presented in the table, the lower the better, best results in bold. Empty cells are due to the impossibility of executing the algorithm for these parameters due to memory problems given the size of the model. Fixed architectures for LAM and Vanilla have been employed, being 7 encoding-decoding layers for LAM and 5 for Vanilla. Count stands for best result counts. Mean Accumulated is computed as the added error divided by the number of valid rows in the table (non-empty). The Vanilla algorithm is the original transformer proposed in [20].
@@
(a) MSE Bayesian test
(b) MAE Bayesian test
Figure 7: Bayesian tests for Vanilla vs LAM
(a) MSE Bayesian test
(b) MAE Bayesian test
Figure 8: Bayesian tests for Informer vs LAM
@@
(a) IHE dataset
(b) NYTaxi dataset
@@
Figure 9: Barplots for the MSE of LAM to showcase the performance depending on the number of encoding-decoding layers used. The datasets shown are IHE and NYTaxi. MSE is presented in the barplot in the y-axis, sequence size (input and output) is presented in the x-axis. The colour of the bar is defined by the number of layers of the model.
@@
(a) RPB dataset
(b) TiNA dataset
@@
Figure 10: Barplots for the MSE of LAM to showcase the performance depending on the number of encoding-decoding layers used. The datasets shown are RPB and TiNa. MSE is presented in the barplot in the y-axis, sequence size (input and output) is presented in the x-axis. The colour of the bar is defined by the number of layers of the model.
@@ 6 Conclusions and future work
This paper proposes a novel attention mechanism for transformers, specifically designed for time series prediction. This mechanism leverages the locality of the data and maintains a theoretical efficiency of $\Theta(log(n))$. Our proposed mechanism has demonstrated superior performance compared to state-of-the-art attention-based mechanisms and models, including Informer, LogFormer, and Reformer, as well as LSTM models.
Additionally, we have highlighted the necessity for new and improved datasets for the evaluation of the Long Sequence Time-series Forecasting (LSTF) problem, in contrast to existing benchmarks. To address this need, we have introduced new datasets that facilitate a more accurate assessment of model quality in these scenarios.
On larger and more challenging datasets, our proposed mechanism LAM has consistently outperformed not only Informer but also the baseline represented by the vanilla model across nearly all scenarios. Furthermore, due to the simplicity of the model accompanying our mechanism, our proposal is uniquely capable of executing effectively in all the studied scenarios. Additionally, we observed that the results exhibit low sensitivity to the depth of the architecture employed. By maintaining a fixed architecture size, we confirmed that the hypothesis remains valid.
@@
Finally, we have identified several areas for improvement and further research to enhance the performance of future models and advance the field. Firstly, as an existing challenge, it is essential to establish a comprehensive benchmark that includes datasets, methodologies, and metrics to evaluate new proposals consistently. Maintaining an up-to-date table of state-of-the-art model results is also crucial. Additionally, our study exclusively proposes an attention mechanism without a dedicated architecture; thus, future work should focus on developing and integrating such architectures. Lastly, considering the promising results obtained with a basic model employing a novel attention mechanism, it is imperative to conduct rigorous ablation studies. These studies will help discern which components positively impact the results and which do not, thereby guiding the development of more effective models.
@@ Acknowledgements
This research results from the Strategic Project IAFER-Cib (C074/23), as a result of the collaboration agreement signed between the National Institute of Cybersecurity (INCIBE) and the University of Granada. This initiative is carried out within the framework of the Recovery, Transformation and Resilience Plan funds, financed by the European Union (Next Generation). Ignacio Aguilera-Martos was supported by the Ministry of Science of Spain under the FPI programme PRE2021-100169.

### v3

@@ 5.1 Experimental results
Informer is a significantly more complex model than the vanilla transformer integrating LAM. The Informer model utilises its own positional encoding based on various time frequencies (seconds, minutes, hours, weeks, etc.) and incorporates layers such as one-dimensional convolutions to facilitate feature extraction from the dataset. Therefore, before proceeding with further experimentation, it is essential to compare LAM and ProbAttn by aligning the frameworks under which both attention mechanisms are implemented. To this end, we will compare the results obtained by ProbAttn and LAM within the vanilla transformer architecture, to isolate the effects of Informer’s attention mechanism with respect to the characteristics of its architecture. This involves using only fully connected layers and the sine-based positional encoding from the original proposal by Vaswani et al (vaswani2017attention, ), instead of the complex architecture of Informer.
@@
The results presented in Table 4 provide a clear illustration of the performance differences. Notably, LAM outperforms ProbAttn in 37 instances, while ProbAttn prevails in 23. However, it is crucial to also examine the cumulative error across all datasets and scenarios to establish a comprehensive metric. Specifically, the cumulative MSE for LAM is 28.97928.97928.979, compared to 34.84134.84134.841 for ProbAttn. This indicates that, overall, LAM exhibits a lower cumulative error than ProbAttn. Similarly, for MAE, LAM achieves an accumulated value of 19.97619.97619.976, whereas ProbAttn records 21.66621.66621.666. Additionally, it is noteworthy that, on average, LAM reduces the error of ProbAttn by approximately 25% for longer prediction horizons. This finding corroborates the improvement and stability of our model’s results under the LSTF problem.
(a) MSE Bayesian test.
(b) MAE Bayesian test.
Figure 6: Bayesian tests ProbAttn vs LAM.
To corroborate the data presented in the table, we include the corresponding Bayesian tests, as illustrated in Figure 6. These tests reveal that the differences are significant in favour of LAM  with a probability of 97.2% for MSE and a probability of 93.9% for MAE of LAM being better than ProbAttn. These results support the effectiveness of LAM within the same comparison framework.
In view of the results obtained in this experimentation, we argue that the complex architecture used in Informer (zhou2023informer, ) is not properly justified. Therefore, the question of determining a transformer architecture that successfully exploits the structure of long-term time series forecasting problems remains open. We note that in the paper that introduced Informer there is no comparison to the vanilla transformer and, thus, no empirical justification of the changes introduced at the architecture level – the study focuses on comparing ProbAttn with FullAttention within their own architecture and comparing Informer to other models specifically designed for time series forecasting. For completeness, we also include comparisons between LAM and FullAttention in Section 6.
@@ 5.2 On the suitability of benchmark models for the LSTF problem
As a conclusion to both experimentations, we have identified several areas for improvement in the literature. Firstly, all datasets utilised have a relatively small number of samples; none exceeds 100,000 instances (Table 1). Transformer models for long-sequence time series forecasting, being highly complex architectures in terms of the number of weights, require a substantial amount of data to effectively train all parameters. The current datasets are insufficient in terms of sample size for this task, which produces results that are not sufficiently conclusive. Additionally, the datasets do not always encompass a large set of features. An increase in the number of variables or features inherently increases the complexity of multivariate prediction for LSTF, leading to more interesting problems.
@@
For these reasons, we believe that the currently used datasets, while useful for measuring model performance, are insufficient to establish a comprehensive benchmark for time series analysis in general, or for the LSTF problem in particular. Consequently, we recognise the necessity of obtaining new datasets that more accurately represent the LSTF problem to better evaluate model performance. This need will be addressed in the following section of this paper.
@@ 6.1 Experimental Details
Datasets: For this experiment, we selected four datasets: Individual Household Electric Power (IHEP) (misc_individual_household_electric_power_consumption_235, ), New York Taxi (NYTaxi) (new_york_taxi, ), Residential Power Battery (RPB) (residential_power_battery, ), and Time-series Industrial Anomaly (TiNA) (tina_dasci_arcelor, ; mhbos, ). The IHEP dataset contains over 200,000 samples with 9 features, representing household electricity consumption measurements over a 4-year period. The NYTaxi dataset includes over 2 million instances with 6 features, such as distance travelled, number of passengers, amount paid, and trip duration. These metrics are aggregated by summing all trips within the same hour, allowing the dataset to support predictions for subsequent hours, days, or weeks. The RPB dataset comprises more than 500,000 samples with 6 features, measuring the actual consumption of a population, solar power generation, and battery levels. Finally, the TiNA dataset includes over 38 million instances with 108 features, sampled every two seconds. This dataset is a real-world dataset provided by ArcelorMittal, reflecting data from an ore mining machine.
@@
Algorithms: In the experimentation provided in Section 4, Informer and LAM are the closest models in terms of performance. The other algorithms, including transformers such as LogTrans and Reformer, showed significantly poorer results. Consequently, we have decided to exclude these algorithms from this comparison to obtain concise and more interpretable results. Therefore, we have retained Informer and LAM  and added the vanilla transformer with its attention mechanism (FullAttention) to serve as a baseline. For our proposed model and the vanilla attention mechanism, the optimal depth was determined through hyperparameter tuning. The transformer architecture employed is a fully connected, symmetric structure with an equal number of encoding and decoding layers. The depth was varied between 2, 3, 5, 7, and 10 layers, contingent upon memory capacity, to identify the best configuration. This parameter will be named as number of encoding-decoding layers for the rest of the section. By using the same architecture to baseline the vanilla attention mechanism, we ensure that the observed results can be solely attributed to our proposal and the network depth. The same standardisation and metrics introduced in the previous section are used for the present one. The train/validation/test split utilised in this study allocated 70% of the data for training, with 15% of this subset further designated for validation and the remaining 85% for training. The remaining 30% of the data was used as the test set.
@@ 6.2 Results and Analysis
Table 6: Multivariate forecasting results, MSE and MAE are presented in the table, the lower the better, best results in bold. Layers represent the best configuration of encoding-decoding layers used to obtain the presented results.
@@
As observed in Table 7, LAM consistently emerges as the superior algorithm in this comparison. Specifically, it outperforms the other algorithms 31 times, whereas Informer obtains better results 5 times, and Vanilla wins 4 times. This demonstrates the clear superiority of our proposal over both the baseline and Informer models. Regarding the mean of the MSE and MAE across rows with valid values, Vanilla appears to have a slight edge. However, it is important to note that these results are significantly influenced by Vanilla’s inability to run under numerous circumstances due to its computational demands. Consequently, these findings strongly establish the validity of our proposal, even when utilising a fixed architecture.
@@
Table 7: Multivariate forecasting results, MSE and MAE are presented in the table, the lower the better, best results in bold. Fixed architectures for LAM and Vanilla have been employed, being 7 encoding-decoding layers for LAM and 5 for Vanilla.
@@
(a) MSE Bayesian test.
(b) MAE Bayesian test.
Figure 7: Bayesian tests for Vanilla vs LAM.
(a) MSE Bayesian test.
(b) MAE Bayesian test.
Figure 8: Bayesian tests for Informer vs LAM.
@@
(a) IHE dataset.
(b) NYTaxi dataset.
@@
Figure 9: Barplots for the MSE of LAMṀSE is presented in the barplot in the y-axis, sequence size (input and output) is presented in the x-axis.
@@
(a) RPB dataset.
(b) TiNA dataset.
@@
Figure 10: Barplots for the MSE of LAMṀSE is presented in the barplot in the y-axis, sequence size (input and output) is presented in the x-axis.
@@ 7 Conclusions and future work
This paper proposes a novel attention mechanism for transformers, specifically designed for long sequence time series forecasting. This mechanism leverages the locality of the data and maintains a theoretical efficiency of $\Theta(log(n))$. Our proposed mechanism has demonstrated superior performance compared to state-of-the-art attention-based mechanisms and models, including Informer (zhou2023informer, ), LogTrans (li2019enhancing, ), Reformer (kitaev2020reformer, ), and Autoformer (wu2021autoformer, ), as well as LSTM models.
Additionally, we have highlighted the necessity for new and improved datasets for the evaluation of the LSTF problem, in contrast to existing benchmarks. To address this need, we have introduced new datasets that facilitate a more accurate assessment of model quality in these scenarios.
On larger and more challenging datasets, LAM has consistently outperformed not only Informer but also the baseline represented by the vanilla model across nearly all scenarios. Furthermore, due to the simplicity of the model accompanying our mechanism, our proposal is uniquely capable of executing effectively in all the studied scenarios. Additionally, we observed that the results exhibit low sensitivity to the depth of the architecture employed. By maintaining a fixed architecture size, we confirmed that the hypothesis remains valid.
@@
Finally, we identified several areas for improvement and further research to enhance the performance of future models and advance the field. Firstly, as an existing challenge, it is essential to establish a comprehensive benchmark that includes datasets, methodologies, and metrics to evaluate new proposals consistently. Maintaining an up-to-date table of state-of-the-art model results is also crucial. Additionally, our study exclusively proposes an attention mechanism without a dedicated architecture; thus, future work should focus on developing and integrating such architectures. Lastly, considering the promising results obtained with a basic model employing a novel attention mechanism, it is imperative to conduct rigorous ablation studies. These studies will help discern which components positively impact the results and which do not, thereby guiding the development of more effective models.
@@ Acknowledgements
This publication is part of the project “Ethical, Responsible, and General Purpose Artificial Intelligence: Applications In Risk Scenarios” (IAFER) Exp.:TSI-100927-2023-1 funded through the creation of university-industry research programs (Enia Programs), aimed at the research and development of artificial intelligence, for its dissemination and education within the framework of the Recovery, Transformation and Resilience Plan from the European Union Next Generation EU through the Ministry of Digital Transformation and the Civil Service. This work was also partially supported by Knowledge Generation Projects, funded by the Spanish Ministry of Science, Innovation, and Universities of Spain under the project PID2023-150070NB-I00. Ignacio Aguilera-Martos was supported by the Ministry of Science of Spain under the FPI programme PRE2021-100169.
Appendix A Theoretical proofs
Proposition
Proof.
We first prove the proposition for $h=1$, that is, we show that
Lemma
Proof.
Lemma
Proof.
$0\leq j-i+L-1=j-i_{1}-rL+L-1=j_{1}-i_{1}\leq j_{1}$
and
$j_{1}$ $=j-(r-1)L-1\leq i-(r-1)L-1=i_{1}+rL-(r-1)L-1$
$=i_{1}+L-1\leq L-1+L-1=2L-2.$
Lemma
Proof.
[…]
  • when $i-L+1\leq j\leq i$, we have $S[i,j]=T_{S}[r,i_{1},j_{1}]$.
This dichotomy concludes the proof. ∎
Lemma
Proof.
[…]
which equals $LAM(Q,K,V)[i,\cdot]$, as we wanted to prove. ∎
Theorem
Proof.
Corollary
Proof.
This follows from instantiating $L$ in Theorem 6. ∎